one man can perform a job in 5 hours
the other can perform the same job in 6 hours
if both work together, what is the time
I am looking for the formula. the answer is 2.73, but I get 2.75

Answer by richwmiller(17219)

x/5+x/6=1
6x+5x=30
11x=30
x=30/11=2.727 abt 2.73
How did you get 2.75?